OpenStudy (anonymous):

Find the linear approximation of the function f(x) = √16-x at a = 0 and use it to approximate the numbers √15.9 and √15.99.... im trying to understand what to do and solving

OpenStudy (anonymous):

To approximate a value of a function at a nasty point (for instance, \[\sqrt{15.99}\]), we can use a nicer value (like 16) and apply an "error" factor. The formula for calculating the linear approximation is given by \[y=f(a) + f \prime(a)(x-a)\] where a is the nice value and x is the nasty value. So, in your case, you have \[f(x) = \sqrt{16-x}, a=0\] \[f\prime(x)=-1/(2 \sqrt{16-x})\] So, \[y=f(a) + f \prime(a)(x-a)\] becomes So, \[y=\sqrt{16-a}-[1/(2 \sqrt{16-a})](x-a)\] Substituting a=0, we get \[y=\sqrt{16}-[1/(2 \sqrt{16})](x)\] \[=4-(x/8).\] So the two values you are looking for correspond to x=.1 and x=.01 (since the original function was 16-x in the radical). Using these values, we get approximations of 3.9875 and3.99875 respectively. \[\] Hope that helps.